OpenStudy (anonymous):

Can someone please help me with this problem x^2= 17/5x+12/5

OpenStudy (campbell_st):

multiply every term by 5x... removes the denominators \[5x^3 = 17 + 12x\] or \[5x^3 - 12x - 17 = 0\]

OpenStudy (anonymous):

Wouldn't you just have to multiply everything by 5 and you would get \[5x^2-17x-12=0\]

OpenStudy (campbell_st):

ok... so the term 12/5x is really \[\frac{12}{5} x ~~~ I ~~thought ~~ \it~~ was ~~~\frac{12}{5x}\] so yes you are correct...

OpenStudy (campbell_st):

if yours is correct you can solve by gruoping in pairs \[5x^2 + 5x - 17x - 17 = 0\] hope it helps